OpenStudy (anonymous):

Evaluate. 35 – 10 ÷ 5 + [(5 + 3) • 4]

OpenStudy (anonymous):

Use The Order Of Operations! PEMDAS.

OpenStudy (anonymous):

Do you have any idea what it may be?

OpenStudy (anonymous):

yes plese my dear aunt sally

OpenStudy (anonymous):

give me madal

OpenStudy (anonymous):

You would change it to: (5+3) (4) 10/5 - 35 = ?

OpenStudy (anonymous):

ol soooooooooooooooooooo

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

I know the answer, but I want to know what you think it is.

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

im not good at this

OpenStudy (anonymous):

how do change pic

OpenStudy (anonymous):

tell me answer to

OpenStudy (anonymous):

Go on settings. I can't tell you the answer, all you need to do is solve this: (5+3) (4) 10/5 - 35 = ? and you'll have your answer!

OpenStudy (anonymous):

Go on settings. I can't tell you the answer, all you need to do is solve this: (5+3) (4) 10/5 - 35 = ? and you'll have your answer!

OpenStudy (anonymous):

Go on settings. I can't tell you the answer, all you need to do is solve this: (5+3) (4) 10/5 - 35 = ? and you'll have your answer!